OpenStudy (anonymous):

what is the vertex of f(x)= x^2-4

OpenStudy (campbell_st):

well for this problem there are several methods 1. completing the square... 2. find the line of symmetry the standard for is \[y = (x -h)^2 + k\] which means the parabola has a vertex at (h, k) so looking at your problem it can be written as \[y = (x - 0)^2 - 4\] compare this to the standard form above and you should be able to see the vertex in the equation.

OpenStudy (campbell_st):

another way to look at it is \[y = x^2 \] has a vertex at (0, 0) what effect does -4 have on the curve... and thats where the vertex moves to.